OpenStudy (anonymous):

Would anyone be able to assist in solving this DE: \[x'' + \frac{ 2U }{ bm } (e ^{-2(x-x _{0})/b} - e ^{(x-x _{0})/b}) = 0\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Sorry, equation dropped off. \[x'' + \frac{ 2U }{ bm } (e ^{-2(x-x _{0})/b} - e ^{(x-x _{0})/b}) = 0\]

OpenStudy (unklerhaukus):

what is U ?, just some constant?

OpenStudy (anonymous):

correct, some constant.

OpenStudy (unklerhaukus):

so it might make it eaisier if you call change 2U/bm = k and change it back when you have solve the equation,

OpenStudy (unklerhaukus):

so that \[x'' + k(e ^{-2(x-x _{0})/b} - e ^{(x-x _{0})/b}) = 0\] now can you see a common factor in the brackets

OpenStudy (anonymous):

good point, thanks Uncle

OpenStudy (anonymous):

I get something that is long and messy as my solution. Can I ask what you get, if you have a moment?

OpenStudy (unklerhaukus):

where did the equation come from , it looks like physics?

OpenStudy (anonymous):

correct, original equation I have is: \[\frac{ 2U }{ b } (e ^{-2(x-x _{0})/b} - e ^{(x-x _{0})/b}) = F(x)\] \[f(x) = mx''\]

OpenStudy (anonymous):

sorry, meant to be \[F(x) = mx''\]

OpenStudy (unklerhaukus):

did you find any common factors in the large brackets?

OpenStudy (anonymous):

you meant the (x - x0)/b?

OpenStudy (unklerhaukus):

yers

OpenStudy (unklerhaukus):

@dumbcow

OpenStudy (anonymous):

Sorry, I may be sleep deprived but I dont see how to proceed. Even if you sub out the the common factor?

OpenStudy (dumbcow):

ok this is a 2nd order non-linear DE which i am not used to and are pretty difficult...i worked it out best i could and i believe there is no closed form "x(t) = ?" solution according to wolfram solution uses hypergeometric function. anyway like @UnkleRhaukus said you can simplify the equation by transforming variables let \[k = \frac{2U}{m}\] \[w = \frac{x-x_0}{b} \rightarrow dw = \frac{dx}{b}\] now DE looks like this \[w'' +\frac{k}{b^{2}}(e^{-2w}-e^{w}) = 0\] now there is a trick to change this to a first order DE.....ref link below http://math.stackexchange.com/questions/42223/solving-the-nonlinear-second-order-ode-u-u2-9-0 this transforms DE to this: \[\frac{1}{2}(w')^{2} = \frac{k}{b^{2}} \int\limits e^{w} -e^{-2w} dw\] after integrating \[\frac{1}{2}(w')^{2} = \frac{k}{b^{2}}(e^{w}+\frac{1}{2}e^{-2w}+c_1)\] solving \[\large w' = \frac{\sqrt{2ke^{w}+ke^{-2w}+c_1}}{b}\] separate variables \[\large \frac{dw}{\sqrt{2ke^{w}+ke^{-2w}+c_1}} = \frac{dt}{b}\] and that integral cant be done using elementary functions.. http://www.wolframalpha.com/input/?i=+dw%2Fsqrt%282e%5Ew%2Be%5E%28-2w%29%29+%3D+dt

OpenStudy (anonymous):

dumbcow, thanks so much for that. I would not have got this...

OpenStudy (unklerhaukus):

hey, great work @dumbcow !

OpenStudy (anonymous):

Yes, thanks both :)